OpenStudy (anonymous):

2(2 - 2x)=-3x a.-4 b.4 c.6 d.0 how do i find the anwser

OpenStudy (anonymous):

thank so much

OpenStudy (nurali):

My pleasure.

OpenStudy (anonymous):

4

OpenStudy (nurali):

2(2 - 2x)=-3x 4-4x=-3x Add 4x on both sides 4-4x+4x=-3x+4x x=4